OpenStudy (anonymous):

Solve the equation sin 5(theta) + sin 3(theta) = 0 please i really need help for this test im going to take in an hour. i really appreciate the help ive already tried different methods.. its not working i hope someone can solve this

OpenStudy (campbell_st):

quick question is it 1. \[\sin^5(\theta) + \cos^3(\theta) = 0\] or 2. \[\sin(5 \theta) + \sin(3\theta)=0\]

OpenStudy (campbell_st):

oops 1. \[\sin^5(\theta) + \sin^3(\theta) = 0\]